# Advantage Alignment Algorithms

### v1

@@ A.1 Deriving the Advantage Alignment Formula
In this section we derive the advantage alignment formula in equation (8) from the opponent shaping expression in equation (6) and assumption 2. Recall assumption 2:
$\pi^{i}(a|s)\propto\exp{Q^{i}(s,a)}$
Recall the opponent shaping policy gradient expression:
@@
$\propto\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta_{1}}\log\exp Q^{2}(s_{t},a_{t},b_{t})\right]$
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta_{1}}Q^{2}(s_{t},a_{t},b_{t})\right]$
For convenience of notation we define:
@@
These are the reward and advantage of agent $i$ at time step $t$ after taking action $a_{t}$ and opponent taking action $b_{t}$. From the Bellman equation (2) we expand as follows:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}Q^{2}(s_{t},a_{t},b_{t})\right]$ (12)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}_{t}\nabla_{\theta^{1}}\left(r^{2}_{t}+\mathbb{E}_{s^{\prime}}\left[V^{2}(s^{\prime})\biggr{|}s_{t},a_{t},b_{t}\right]\right)\right]$ (13)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}_{t}\nabla_{\theta^{1}}\mathbb{E}_{s^{\prime}}\left[V^{2}(s^{\prime})\biggr{|}s_{t},a_{t},b_{t}\right]\right]$ (14)
$=\mathbb{E}\underset{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}{}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}_{t}\mathbb{E}_{\tau^{\prime}\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{k=0}^{\infty}\gamma^{k}A^{2}_{k}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}|s_{k})\biggr{|}s_{t},a_{t},b_{t}\right]\right]$ (15)
$=\mathbb{E}\underset{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}{}\left[\sum_{t=0}^{\infty}\mathbb{E}_{\tau^{\prime}\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{k=0}^{\infty}\gamma^{k+t}A^{1}_{t}A^{2}_{k}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}|s_{k})\biggr{|}s_{t},a_{t},b_{t}\right]\right]$ (16)
$=\mathbb{E}\underset{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}{}\left[\mathbb{E}_{\tau^{\prime}\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}\gamma^{k}A^{1}_{t}A^{2}_{k}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}|s_{k})\biggr{|}s_{t},a_{t},b_{t}\right]\right]$ (17)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}\gamma^{k}A^{1}_{t}A^{2}_{k}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}|s_{k})\right]$ (18)
We use the Bellman equation in line (13). In line (16), we use the fact that $\gamma^{t}A^{1}(s_{t},a_{t},b_{t})$ is measurable w.r.t. the natural filtration of the process up to time $t$, $\mathcal{F}_{t}$, and the independence of the two terms conditioned on $\mathcal{F}_{t}$ by the Markov property. In line (17) we use linearity of expectation. In line (18) we use the rule of the iterated expectations.
@@ A.2 Proof of Theorem 1
LOLA (Foerster et al., 2018b, ) optimizes the return of the agent under an imagined optimization step of the opponent (assuming the opponent is a naive learning algorithm). Under their notation, a LOLA agent optimizes $V^{1}(\theta_{1},\theta_{2}+\Delta\theta_{2})$ where $\Delta\theta_{2}$ is a gradient ascent step on the parameters of the opponent $\theta_{2}$. Since computing this value function explicitly is difficult, LOLA uses the first-order Taylor expansion surrogate objective:
@@
$\nabla_{\theta_{1}}V^{1}(\theta^{1},\theta^{2}+\nabla_{\theta_{2}}V^{2}(\theta_{1},\theta_{2}))=\nabla_{\theta_{1}}V^{1}(\theta_{1},\theta_{2})+\left(\nabla_{\theta_{1}}\nabla_{\theta_{2}}V^{2}(\theta^{1},\theta^{2})\right)\nabla_{\theta_{2}}V^{1}(\theta_{1},\theta_{2}).$ (20)
@@
The first-order terms above can be computed using the Advantage form of the REINFORCE estimator, which is given by equation (3). Foerster et al., 2018b derive the following REINFORCE estimator for the second-order term:
@@
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r^{2}_{t}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{k}|s_{k})\right)\left(\sum_{k=0}^{t}\nabla_{\theta_{2}}\log\pi^{2}(b_{k}|s_{k})\right)^{T}\right].$ (21)
For convenience of notation we construct the time dependent sets:
$B_{t}:=\{\nabla_{\theta_{1}}\log\pi^{1}(a_{0}|s_{0}),\ldots,\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})\}$
$C_{t}:=\{\nabla_{\theta_{2}}\log\pi^{2}(b_{0}|s_{0}),\ldots,\>\nabla_{\theta_{2}}\log\pi^{2}(b_{t}|s_{t})\}$
@@
Using this notation, we expand the second order term beginning from equation (20), to bring out the advantage $A_{t}^{2}$:
$\nabla_{\theta_{1}}\nabla_{\theta_{2}}V^{2}(\theta^{1},\theta^{2})$ (22)
@@
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r^{2}_{t}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})\right)\left(\sum_{k=0}^{t}\nabla_{\theta_{2}}\log\pi^{2}(b_{t}|s_{t})\right)^{T}\right]$ (23)
@@
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{\nabla_{a}\in B_{t}}\sum_{\nabla_{b}\in C_{t}}\nabla_{a}\nabla_{b}^{T}\right)\sum_{k=t}^{\infty}\gamma^{k}r^{2}_{k}\right]$ (24)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{\nabla_{a}\in B_{t}}\sum_{\nabla_{b}\in C_{t}}\nabla_{a}\nabla_{b}^{T}\right)R^{2}(s_{t})\right]$ (25)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{\nabla_{a}\in B_{t}}\sum_{\nabla_{b}\in C_{t}}\nabla_{a}\nabla_{b}^{T}\right)\mathbb{E}\left[R^{2}(s_{t})|s_{t},a_{t}\right]\right]$ (26)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{\nabla_{a}\in B_{t}}\sum_{\nabla_{b}\in C_{t}}\nabla_{a}\nabla_{b}^{T}\right)Q^{2}(s_{t},b_{t})\right]$ (27)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{\nabla_{a}\in B_{t}}\sum_{\nabla_{b}\in C_{t}}\nabla_{a}\nabla_{b}^{T}\right)A^{2}_{t}\right]$ (28)
Where we reorder the terms of the summation to sum over future rewards instead of past gradient terms in line 24, we use the law of iterated expectation in line 26 and a baseline subtraction in line 28. $R^{2}(s_{t})$ denotes the return starting at state $s_{t}$.
Per Equation (20), we multiply this Hessian with the gradient $\nabla_{\theta_{2}}V^{1}(\theta_{1},\theta_{2})$ to obtain
@@
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}A^{2}_{t}\left(\sum_{\nabla_{a}\in B_{t}}\sum_{\nabla_{b}\in C_{t}}\nabla_{a}\nabla_{b}^{T}\right)\right]\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}A^{1}_{t}\nabla_{\theta_{2}}\log\pi^{2}(b_{t}|s_{t})\right].$
This completes the proof. ∎

### v3

@@ A.1 Deriving the Advantage Alignment Formula
In this section we derive the Advantage Alignment formula in equation equation 8 from the opponent shaping expression in equation equation 6 and assumption 2. Recall assumption 2:
$\pi^{i}(a|s)\propto\exp{\beta\mathbb{E}_{b\sim\pi^{3-i}(\cdot|s)}[Q^{i}(s,a,b)]}$
Note that if $i=1,3-i=2$ and if $i=2,3-i=1$. Recall the opponent shaping policy gradient expression:
@@
$=\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta_{1}}\log\exp\mathbb{E}_{a\sim\pi^{1}(|s_{t})}[Q^{2}(s_{t},a,b_{t})]\right]$
$=\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta_{1}}\mathbb{E}_{a\sim\pi^{1}(|s_{t})}[Q^{2}(s_{t},a,b_{t})]\right].$
where is the second line we used the fact that the expected advantage is zero. For convenience of notation we define:
@@
These are the reward and advantage of agent $i$ at time step $t$ after taking action $a_{t}$ and opponent taking action $b_{t}$. From the Bellman equation equation 2 we expand as follows:
$\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}\mathbb{E}_{a\sim\pi^{1}(|s_{t})}[Q^{2}(s_{t},a,b_{t})]\right]$ (13)
$=\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}_{t}\nabla_{\theta^{1}}\left(\mathbb{E}_{s^{\prime}}\left[r^{2}_{t}+\gamma\cdot V^{2}(s^{\prime})\biggr{|}s_{t},b_{t}\right]\right)\right]$ (14)
$=\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t+1}A^{1}_{t}\nabla_{\theta^{1}}\mathbb{E}_{s^{\prime}}\left[V^{2}(s^{\prime})\biggr{|}s_{t},b_{t}\right]\right]$ (15)
$=\beta\cdot\mathbb{E}\underset{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}{}\left[\sum_{t=0}^{\infty}\gamma^{t+1}A^{1}_{t}\mathbb{E}_{\tau^{\prime}\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{k=0}^{\infty}\gamma^{k}A^{2}_{k}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}^{\prime}|s_{k}^{\prime})\biggr{|}s_{t},b_{t}\right]\right]$ (16)
$=\beta\cdot\mathbb{E}\underset{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}{}\left[\sum_{t=0}^{\infty}\mathbb{E}_{\tau^{\prime}\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{k=0}^{\infty}\gamma^{k+t+1}A^{1}_{t}A^{2}_{k}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}^{\prime}|s_{k}^{\prime})\biggr{|}s_{t},b_{t}\right]\right]$ (17)
$=\beta\cdot\mathbb{E}\underset{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}{}\left[\mathbb{E}_{\tau^{\prime}\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}\gamma^{t+k+1}A^{1}_{t}A^{2}_{k}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}^{\prime}|s_{k}^{\prime})\biggr{|}s_{t},b_{t}\right]\right]$ (18)
$=\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}\gamma^{t+k+1}A^{1}_{t}A^{2}_{k}\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}|s_{k})\right].$ (19)
We use Bellman equation in line (14), and the policy gradient theorem in line (16). In line (17), we use the fact that $\gamma^{t}A^{1}(s_{t},a_{t},b_{t})$ is measurable w.r.t. the natural filtration of the process up to time $t$, $\mathcal{F}_{t}$, and the independence of the two terms conditioned on $\mathcal{F}_{t}$ by the Markov property. In line (18) we use linearity of expectation. In line (19) we use the rule of the iterated expectations. Reorganizing the summations in causal form we get the desired result:
$\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t+1}\left(\sum_{k<t}\gamma^{t-k}A^{1}(s_{k},a_{k},b_{k})\right)A^{2}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{t}|s_{t})\right].$ (8)
@@ A.2 Proof of Theorem 1
Lemma 1 (Policy changes under gradient ascent).
Given a policy $\pi_{\theta}(a|s)$ parametrised such that the set of gradients $\nabla_{\theta}\log\pi_{\theta}(a|s)$ for all pairs $(a,s)$ form an orthonormal basis, and a value function $V(\theta)$, the following holds:
$\frac{d}{d\alpha}\pi_{\theta+\alpha\nabla_{\theta}V}(a|s)=\nabla_{\theta}V\cdot\nabla_{\theta}\pi_{\theta}(a|s)$ $=d_{\gamma}(s)\pi_{\theta}(a|s)^{2}A(a|s)$ (20)
$d_{\gamma}(s)$ $\equiv\sum_{t=0}^{\infty}\gamma^{t}Pr(s_{t}=s)$ (21)
$A(a|s)$ $\equiv Q(a|s)-V(s)$ (22)
Proof.
The policy gradient theorem gives us:
$\nabla_{\theta}V(\theta)$ $=\mathbb{E}_{\tau\sim\pi}\sum_{t=0}^{\infty}\gamma^{t}A(a_{t}|s_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})$ (23)
$=\sum_{(s,a)}d_{\gamma}(s)\pi_{\theta}(a|s)A(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)$ (24)
Taking the dot product of this expression with $\nabla_{\theta}\pi_{\theta}(a^{\prime}|s^{\prime})=\pi_{\theta}(a^{\prime}|s^{\prime})\nabla_{\theta}\log\pi_{\theta}(a^{\prime}|s^{\prime})$ and invoking the assumed orthonormality of gradients:
$\nabla_{\theta}\pi_{\theta}(a^{\prime}|s^{\prime})\cdot\nabla_{\theta}V(\theta)$ (25)
$=\sum_{(s,a)}d_{\gamma}(s)\pi_{\theta}(a|s)\pi_{\theta}(a^{\prime}|s^{\prime})A(a|s)\bigg{(}\nabla_{\theta}\log\pi_{\theta}(a|s)\cdot\nabla_{\theta}\log\pi_{\theta}(a^{\prime}|s^{\prime})\bigg{)}$ (26)
$=d_{\gamma}(s^{\prime})\pi_{\theta}(a^{\prime}|s^{\prime})^{2}A(a^{\prime}|s^{\prime})$ (27)
Theorem 1. (LOLA policy gradient) Given a two-player game where players 1 and 2 have respective policies $\pi^{1}(a|s)$ and $\pi^{2}(b|s)$, where each policy is parametrised such that the set of gradients $\nabla_{\theta_{2}}\log\pi^{2}(a|s)$ for all pairs $(a,s)$ form an orthonormal basis, the LOLA update for the first player correspond to a reinforce update with an advantage
$A_{LOLA}^{*}(s_{t},a_{t},b_{t})=A^{1}(s_{t},a_{t},b_{t})+\beta\cdot\sum_{k=t}^{\infty}d_{\gamma,{k-t}}\gamma^{k-t}A^{1}_{k}A^{2}_{k-t}.$ (11)
where $A^{i}_{k}:=A^{i}(s_{k},a_{k},b_{k})$ and $d_{\gamma,k}$ is the occupancy measure of the tuple $(a_{k},b_{k},s_{k})$
Proof.
LOLA (Foerster et al., 2018b, ) optimizes the return of the agent under an imagined optimization step of the opponent (assuming the opponent is a naive learning algorithm). Under their notation, a LOLA agent optimizes $V^{1}(\theta_{1},\theta_{2}+\Delta\theta_{2})$ where $\Delta\theta_{2}$ is a gradient ascent step on the parameters of the opponent $\theta_{2}$. Note that along this proof because we consider the method proposed by (Foerster et al., 2018b, ) we use their way to compute gradients. Particularly, one does not use Assumption 2, and consequently assume that $\nabla_{\theta_{1}}\log\pi_{\theta_{2}}=0$ (and respectively $\nabla_{\theta_{2}}\log\pi_{\theta_{1}}=0$.)
Since computing this value function explicitly is difficult, LOLA uses the first-order Taylor expansion surrogate objective:
@@
$\nabla_{\theta_{1}}V^{1}(\theta^{1},\theta^{2}+\alpha\nabla_{\theta_{2}}V^{2}(\theta_{1},\theta_{2}))=\nabla_{\theta_{1}}V^{1}(\theta_{1},\theta_{2})+\beta\left(\nabla_{\theta_{1}}\nabla_{\theta_{2}}V^{1}(\theta^{1},\theta^{2})\right)\nabla_{\theta_{2}}V^{2}(\theta_{1},\theta_{2}).$ (29)
@@
The first-order terms above is computed using the Advantage form of the REINFORCE estimator, which is given by equation equation 3. Foerster et al., 2018b derive the following REINFORCE estimator for the second-order term:
$\nabla_{\theta_{1}}\nabla_{\theta_{2}}V^{1}(\theta^{1},\theta^{2})$ (30)
@@
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r^{1}_{t}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})\right)\left(\sum_{k=0}^{t}\nabla_{\theta_{2}}\log\pi^{2}(b_{t}|s_{t})\right)^{\top}\right]$ (31)
Now, we use the following fact
$\sum_{t=0}^{\infty}c_{t}(\sum_{k=0}^{t}a_{k})(\sum_{l=0}^{t}b_{l})=\sum_{t=0}^{\infty}c_{t}\sum_{S=0}^{t}\sum_{k=0}^{S}a_{k}b_{S-k}=\sum_{S=0}^{\infty}\sum_{t=S}^{\infty}\sum_{k=0}^{S}a_{k}b_{S-k}c_{t}=\sum_{S=0}^{\infty}\sum_{k=0}^{S}a_{k}b_{S-k}\sum_{t=S}^{\infty}c_{t}$
@@
to expand the second order term beginning from equation 29, to bring out the advantage $A_{t}^{1}$:
$\nabla_{\theta_{1}}\nabla_{\theta_{2}}V^{1}(\theta^{1},\theta^{2})$ (32)
@@
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r^{1}_{t}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})\right)\left(\sum_{k=0}^{t}\nabla_{\theta_{2}}\log\pi^{2}(b_{t}|s_{t})\right)^{\top}\right]$ (33)
@@
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{S=0}^{\infty}\left(\sum_{k=0}^{S}\nabla_{\theta_{1}}\log\pi^{1}(a_{k}|s_{k})\nabla_{\theta_{2}}\log\pi^{2}(b_{S-k}|s_{S-k})^{\top}\right)\sum_{t=S}^{\infty}\gamma^{t}r^{1}_{t}\right]$ (34)
@@
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{k}|s_{k})\nabla_{\theta_{2}}\log\pi^{2}(b_{t-k}|s_{t-k})^{\top}\right)\sum_{l=t}^{\infty}\gamma^{l}r^{1}_{l}\right]$ (35)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{k}|s_{k})\nabla_{\theta_{2}}\log\pi^{2}(b_{t-k}|s_{t-k})^{\top}\right)\gamma^{t}\mathbb{E}\left[\sum_{l=0}^{\infty}\gamma^{l}r^{1}_{l+t}\right]\right]$ (36)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{k}|s_{k})\nabla_{\theta_{2}}\log\pi^{2}(b_{t-k}|s_{t-k})^{\top}\right)\gamma^{t}Q^{1}(s_{t},a_{t},b_{t})\right]$ (37)
$=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{k}|s_{k})\nabla_{\theta_{2}}\log\pi^{2}(b_{t-k}|s_{t-k})^{\top}\right)\gamma^{t}A^{1}_{t}\right],$ (38)
where we reorder the terms of the summation to sum over future rewards instead of past gradient terms in line 34, we use the law of iterated expectation in line 36 and a baseline subtraction in line 38.
Per Equation equation 29, we multiply this Hessian with the gradient of the value function
$\nabla_{\theta_{2}}V^{2}(\theta_{1},\theta_{2})$ $=\mathbb{E}_{\tau\sim\pi}\sum_{t=0}^{\infty}\gamma^{t}A^{2}(a_{t},b_{t}|s_{t})\nabla_{\theta_{2}}\log\pi^{2}(b_{t}|s_{t})$ (39)
$=\sum_{(s,a,b)}d_{\gamma}(a,b,s)A^{2}(a,b|s)\nabla_{\theta_{2}}\log\pi^{2}(b|s)$ (40)
where where $d_{\gamma}(a,b,s)$ is the occupancy measure of the state actions tuple $(a,b,s)$, and use the assumption that the gradients $(\nabla_{\theta_{2}}\log\pi^{2}(a|s))_{(a,s)}$ form an orthonormal basis to obtain
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{k=0}^{t}\nabla_{\theta_{1}}\log\pi^{1}(a_{k}|s_{k})d_{\gamma}(a_{t-k},b_{t-k},s_{t-k})A^{2}_{t-k}\right)\gamma^{t}A^{1}_{t}\right].$
To completes the proof, we finally need to switch the summations to get
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\left(\sum_{k=t}^{\infty}\gamma^{k}A^{1}_{k}d_{\gamma}(a_{k-t},b_{k-t},s_{k-t})A^{2}_{k-t}\right)\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})\right].$